OpenStudy (anonymous):

Express the following as a partial fraction: (i) 4x+5/(x+2)^2

OpenStudy (anonymous):

4x + 5 = 4/(x+2) - 3/(x+2)^2

OpenStudy (anonymous):

A=4 , and B= -3

OpenStudy (anonymous):

hina your right how did u get it ?

OpenStudy (anonymous):

(4x + 5)/(x+2)^2 = A/(x+2) + B/(x+2)^2

OpenStudy (anonymous):

then we multiply n divide both sides by LCM i.e. (x+2)^2 to get the value of B, we need to put x+2 = 0, which gives x=-2 then B= -3 for getting value of A, we have, 4x+5 = A(x+2)-3 opeing bracket 4x+5 = Ax + 2A-3 then compare coefficients and you will get the required answer

OpenStudy (anonymous):

ty:)